OpenStudy (anonymous):

Solve the following system using the substitution method: 6x - 2y = -4 y = 3x + 2

OpenStudy (anonymous):

but how would I go about it in this case?

OpenStudy (anonymous):

what would the solution be like?

hero (hero):

In this case, substitute the expression 3x + 2 into the first equation to get: 6x - 2(3x + 2) = -4 Then simplify and isolate x. Afterwards, find the value of y.

OpenStudy (anonymous):

thx but look here: http://www.wolframalpha.com/input/?i=6x+-+2%283x+%2B+2%29+%3D+-4

OpenStudy (anonymous):

is it Inconsistent? Dependent? or are the coordinates? what would the answer be like? thx so much!

hero (hero):

Actually, the first thing I usually do before using any method is to simplify equations when necessary. We can simplify the first equation. When we do that we get: 3x - y = 2 That can be manipulated to get 3x + 2 = y Which so happens to be the same line as the second equation. So no need to use substitution method since both equations are actually the same.

OpenStudy (anonymous):

so which coordinate would that be? or would it be Inconsistent or Dependent? thx by the way for your time. :D

OpenStudy (callisto):

Keywords: "since both equations are actually the same. "

OpenStudy (anonymous):

so Inconsistent or Dependent?

OpenStudy (callisto):

Inconsistent is like... x+y = 1 and x + y =5 Dependent is like... x+y = 1 2x + 2y= 2 Any guess?

OpenStudy (anonymous):

Nope I'm bricked... :( Can you PLEASE provide the response! :(

OpenStudy (dumbsearch2):

what's the graphing lines? what's the coordinates? or is it dependent? or inconsistent?

OpenStudy (dumbsearch2):

Look here: